OpenStudy (texaschic101):

I need to borrow somebody's brain for a second...just need help with the set-up...not the answer

OpenStudy (texaschic101):

Peter has 48 coins. He has three times as many 10-cent coins as 20-cent coins and the remainder are 5-cent coins. Determine the number of each coin type if their total value is $5.10 x = 10 cent coins, y = 20 cent coins and z = 5 cent coins I have this, but I think it is wrong x + y + z = 48 3y = x x - y = z .10x + .20y + .05z = 5.10 but my answers are not coming out right

OpenStudy (texaschic101):

I am getting confused on " remainder is 5 cent coins "

jimthompson5910 (jim_thompson5910):

I'm not sure why they put that part in there, but you can get away with not using "x - y = z" (not sure how you got that equation?) you just need 3 equations to solve for 3 variables. So you can use x + y + z = 48 3y = x .10x + .20y + .05z = 5.10

OpenStudy (texaschic101):

But what about the remainder being 5 cent coins.....don't need that ?

OpenStudy (texaschic101):

because.. 3y + y + z = 48 4y + z = 48 4y = 48 - z y = 12 - z/4

jimthompson5910 (jim_thompson5910):

I'm not sure why they wrote that. It's awkwardly phrased and unneeded.

jimthompson5910 (jim_thompson5910):

It might be better to solve 4y + z = 48 for z (to avoid getting fractions)

OpenStudy (texaschic101):

4y + z = 48 z = 48 - 4y

jimthompson5910 (jim_thompson5910):

.10x + .20y + .05z = 5.10 .10x + .20y + .05(48-4y) = 5.10 ... plug in z = 48 - 4y .10(3y) + .20y + .05(48-4y) = 5.10 ... plug in x = 3y solve for y

OpenStudy (texaschic101):

.10(3y) + .20y + .05(48 - 4y) = 5.10 .30y + .20y + 2.40 - .20y = 5.10 .30y = 5.10 - 2.40 .30y = 2.70 y = 9 3y = x 3(9) = x 36 = x x + y + z = 48 36 + 9 + z = 48 45 + z = 48 z = 3 that works

OpenStudy (texaschic101):

thank you soooo much. I got lost on this one

jimthompson5910 (jim_thompson5910):

you're welcome

jimthompson5910 (jim_thompson5910):

y = 9 is correct however, x = 36 and z = 3 are incorrect

OpenStudy (texaschic101):

awww....I guess I should have checked it. I just saw that there was 48 coins

OpenStudy (texaschic101):

27 = x....stupid mistake...I can get it now

OpenStudy (texaschic101):

your a life-saver

jimthompson5910 (jim_thompson5910):

glad to be of help